OpenStudy (anonymous):

How do I prove that 2 cot 2x = cot x-tan x? Thank you! I'd appreciate it!

OpenStudy (asnaseer):

use the identity:\[\cot(2x)=\frac{\cot^2(x)-1}{2\cot(x)}\]

OpenStudy (anonymous):

Where did you get cot\[^{2}\] at?

OpenStudy (anonymous):

cot^2

OpenStudy (asnaseer):

it comes from the re-arrangement of:\[\tan(2x)=\frac{2\tan(x)}{1-\tan^2(x)}\]

OpenStudy (asnaseer):

you know:\[\cot(x)=\frac{1}{\tan(x)}\]

OpenStudy (anonymous):

In fact, just take the reciprocal of the last given identity...

OpenStudy (anonymous):

The tan one, the answer drops out...

OpenStudy (asnaseer):

jennifer473: are you aware of these identities? you need to know them in order to prove your question.

OpenStudy (anonymous):

Yes I know the identities. It's just hard for me to really see where to put the identity. I know that I should only work on one side to match the other. But when I tried working through this problem on my exam a few days ago, I did not arrive at the answer. I kept going around in circles.

OpenStudy (asnaseer):

ok, so if we start from:\[\tan(2x)=\frac{2\tan(x)}{1-\tan^2(x)}\]and invert this identity, we get:\[\frac{1}{\tan(2x)}=\frac{1-\tan^2(x)}{2\tan(x)}\]follow so far?

OpenStudy (anonymous):

I am able to do that even though the problem starts with 2cot2x?

OpenStudy (asnaseer):

yes - just see if you can follow along - it will all become clear soon...

OpenStudy (asnaseer):

so next, we know:\[\frac{1}{\tan(2x)}=\cot(2x)\]agreed?

OpenStudy (anonymous):

Correct

OpenStudy (asnaseer):

so we substitute this into my last result to get:\[\frac{1}{\tan(2x)}=\frac{1-\tan^2(x)}{2\tan(x)}\]therefore:\[\cot(2x)=\frac{1-\tan^2(x)}{2\tan(x)}\]now multiply both sides by 2 to get:\[2\cot(2x)=\frac{1-\tan^2(x)}{2\tan(x)}\]make sense?

OpenStudy (asnaseer):

sorry, last line should be:\[2\cot(2x)=\frac{1-\tan^2(x)}{\tan(x)}\]

OpenStudy (asnaseer):

making sense?

OpenStudy (anonymous):

We are working on just the left side of the equation correct?

OpenStudy (asnaseer):

we started with an identity for tan(2x) and are re-arranging it to get the desired result.

OpenStudy (anonymous):

Oh I see. It's making a little sense to me.

OpenStudy (asnaseer):

ok, so final step is:\[2\cot(2x)=\frac{1-\tan^2(x)}{\tan(x)}=\frac{1}{\tan(x)}-\frac{\tan^2(x)}{\tan(x)}=\cot(x)-\tan(x)\]

OpenStudy (asnaseer):

I hope my explanation made sense

OpenStudy (anonymous):

So for 2cot2x, I bring the 2 from the 2x to the front? What happened to the 2 that was in front of 2cot2x?

OpenStudy (asnaseer):

sorry - I don't understand your question?

OpenStudy (anonymous):

The original problem is to prove that 2cot2x = cotx-tanx. So what happened to the "2" that is in front of cot2x?

OpenStudy (asnaseer):

look at the last line of my result - it is there

OpenStudy (asnaseer):

we started with the identity:\[\tan(2x)=\frac{2\tan(x)}{1-\tan^2(x)}\]which we inverted to get:\[\frac{1}{\tan(2x)}=\frac{1-\tan^2(x)}{2\tan(x)}\]and then used the fact that \(\frac{1}{\tan(2x)}=\cot(2x)\) to get:\[\cot(2x)=\frac{1-\tan^2(x)}{2\tan(x)}\]then multiplied both sides by 2 to get:\[2\cot(2x)=\frac{1-\tan^2(x)}{\tan(x)}=\cot(x)-\tan(x)\]

OpenStudy (anonymous):

Oh! I see now! I feel dumb for not noticing that earlier. Lol. Thank you so much. I really appreciate you taking the time out to explain this step by step. It has really helped me understand how I got it wrong on the exam. So thank you!

OpenStudy (asnaseer):

no problem - we all have to start somewhere - so don't feel as if you are "dumb" - you definitely are not. :)

OpenStudy (anonymous):

Thank you :) Have a lovely rest of the day!

OpenStudy (asnaseer):

I will thanks - you too :)